# Robustness Analysis for Quantum Systems Controlled by Continuous-Time Pulses

### v2

@@ Abstract
I Introduction
@@
We contribute to the ongoing development of robust quantum control by presenting a procedure to quantify robustness for continuously-varying, smooth controls and applicable to multi-level systems. Although piecewise constant controls have become a dominant paradigm in quantum control, such controls yield “solutions” of the Schrödinger equation that are piecewise defined and continuous but not differentiable at the switching times. Although further work is needed to explore the advantages of continuously-varying and smooth controls, there are theoretical and practical reasons why smooth controls may be preferable for certain applications. Extending our framework to evaluate robustness to such controls is therefore important. In Section II, we generalize the sensitivity analysis of [6] and [7] to accommodate dynamics governed by continuously varying controls. As part of this analysis, we introduce the Magnus expansion as a critical tool in the analysis of systems controlled by pulses that depart from the piecewise constant paradigm. After providing conditions for vanishing sensitivity (optimal robustness) in Section III, we develop a bound for the sensitivity to any physically-realizable and norm-bounded uncertainty in Section IV. Finally, Section V presents two quantum technology-inspired case studies, based on state transfer in a spintronic network and the implementation of a quantum NOT gate.
II Preliminaries
II-A System Dynamics and the Magnus Expansion
@@
II-B Approximate Solutions Given by the Magnus Expansion
@@
with $t_{k}=kh$ for $k\in\set{0,1,\ldots\tau-1}$. This is similar to the solution for piecewise constant controls, and the general approximation for the propagator over the interval $t_{\tau_{1}},t_{\tau_{2}})$ is
@@
Given the complexity of the higher order terms in the Magnus expansion and intractability of providing closed-form expressions for the $\Omega_{j}(t)$, typical application of the Magnus approach requires truncation of the series at a given order. Numerous methods of various orders have been derived [9]. For the numerical examples in Section V, we adopt a fourth-order method to approximate the propagator over the time interval $t_{k-1},t_{k})$ for each $k$ from $1$ to $\tau$ as $U^{(k,k-1)}:=\exp(\Gamma^{(k)})$. Following [11, 12] the approximate dynamical generator is defined as $\Gamma^{(k)}:=$
@@
where $A(t_{k})=A^{(k)}$, $A(t_{k}+h/2)={A}^{(k+1/2)}$. As this method is fourth-order, the error of the exact propagator denoted $\mathcal{U}^{(k,k-1)}$ over the time-step (k1)h(k-1)h to $kh$ compared to the truncated estimate, $\|\mathcal{U}^{(k,k-1)}-U^{(k,k-1)}\|$, is $\mathcal{O}(h^{5})$ [9, 11, 13, 14, 15].
II-C Fidelity and Performance
@@
II-D First-Order Sensitivity of the Performance
@@
$X_{\mu}^{(k)}=U^{(\tau,k)}\left(\partial_{\mu}U^{(k,k-1)}\right)U^{(k-1,0)}.$ (12)
@@ III Vanishing Sensitivity
[19, Th. 21.7] If the fidelity map $\mathsf{F}:\mathfrak{u}(N)\rightarrow[0,1]$ is differentiable except possibly at some isolated points, then $\forall U\in\mathsf{F}^{-1}(1)$ where the differential exists, $d_{U}\mathsf{F}=0$.
Viewing the physically relevant, skew-hermitian $S_{\mu}$ as providing the direction along which the sensitivity is computed, this sensitivity is $d_{{A}}\mathsf{F}(S_{\mu})$ as seen from (15).
@@
IV Bounding the Differential Sensitivity to Arbitrary Perturbations
At the other end of the spectrum, we consider upper bounds on the potential size the differential sensitivity for a given controller for a set of allowable uncertainties or perturbations to the dynamics. To better analyze the relationship between a set of possible perturbation structures, represented by the family $\set{\partial_{\mu}\Gamma^{(k)}}$, and the resulting sensitivity, we seek an expression such that the perturbation structure encoded in $\partial_{\mu}\Gamma^{(k)}$ may be viewed as the input to a controller-specific operator that yields the resulting sensitivity to a given structure. Central to this derivation is the controller-specific mapping
@@ IV Bounding the Differential Sensitivity to Arbitrary Perturbations
Proof:
For brevity write $\mathcal{D}^{(k)}\left(\partial_{\mu}\Gamma^{(k)}\right)$ as $X$ and $P^{(k)}z^{*}$ as $Y$. By the properties of the adjoint action of a Lie group element on its Lie algebra, $X$ is in $\mathfrak{u}(N)$ and so skew-Hermitian. Decompose $Y$ as $Y=Y_{H}+Y_{SH}$ where $Y_{H}$ is Hermitian and $Y_{SH}$ is skew-Hermitian. It then follows that $\Re\left\{\operatorname{\mathrm{Tr}}\left(XY\right)\right\}=\Re\left\{\operatorname{\mathrm{Tr}}\left(XY_{SH}\right)\right\}$ as $\Re\left\{\operatorname{\mathrm{Tr}}\left(XY_{H}\right)\right\}=0$. ∎
@@
Proof:
The proof follows from the exposition above, which gives $\partial_{\mu}\tilde{\mathsf{F}}_{\alpha}=\gamma\sum_{k=1}^{\tau}\mathsf{z}^{(k)}\mathsf{g}_{\mu}^{(k)}$. Taking absolute values on both sides of the equation and employing the triangle equality to the sum yields $|\partial_{\mu}\mathsf{F}_{\alpha}|\leq\gamma\sum_{k=1}^{\tau}\left|\mathsf{z}^{(k)}\mathsf{g}_{\mu}^{(k)}\right|$. Applying Cauchy-Schwarz to the inner product of $\mathsf{z}^{(k)}$ with $\mathsf{g}_{\mu}^{(k)}$ for each $k$ and enforcing the bound on $\|\mathsf{g}_{\mu}^{(k)}\|\leq b$ we have $\left|\partial_{\mu}\tilde{\mathsf{F}}_{\alpha}\right|\leq\gamma\left(\sum_{k=1}^{\tau}\|\mathsf{z}^{(k)}\|\right)b=\gamma\|\mathbf{Z}\|_{1}b$. ∎
This construction allows us to consider the controller and input-output state information encoded in sequence $\mathsf{z}^{(k)}$ separately from the perturbation-specific effects encoded in $\partial_{\mu}\Gamma^{(k)}$, and efficiently evaluate the worst-case sensitivity of a given controller for a collection of perturbations $\set{\partial_{\mu}\Gamma^{(k)}}_{\mu}$. Instead of bounding the sensitivity of the fidelity for a given controller with regard to certain perturbations, one can alternatively derive lower bounds on the fidelity considering a range of perturbations, an approach pursued in [21, Th. 1]. Both approaches can be seen as complementary.
V Applications
We now apply results of Section IV to two different quantum technology platforms subject to distinct time-varying control protocols. We do this to illustrate that the results are not dependent upon a specific application or control pulses synthesized from a specific set of basis functions.
V-A State Transfer in Spin Ring
@@
Figure 1: Plot comparing $|\partial_{\mu}\mathsf{F}_{S}|$ for $\mu=\set{0,1,2}$ with $\beta$ from Theorem 2. Though computed only from the nominal system and control data, the bound has an accuracy to within slightly more than one order of magnitude. The controllers are ordered by increasing fidelity so that the controller with index 1 yields the smallest fidelity and that with index 8 yields the greatest.
Given a set of high fidelity controllers, the differential sensitivity to collective uncertainty in any of the Hamiltonian matrices is readily computed by application of (10) and (12) to (17). Based on the bound $\mathcal{B}$ used in the synthesis, we compute the maximum norm of the perturbations $\partial_{\mu}\Gamma^{(k)}$ as $b=\max\limits_{\mu,k}\|\partial_{\mu}\Gamma^{(k)}\|=0.1555$. We then test the bound $\beta$ of Theorem 2 after computing $\|\mathbf{Z}\|_{1}$ as derived in Section IV. Figure 1 shows the result. Without the need to explicitly compute the derivative of a matrix exponential as in (13), and based simply on knowledge of the controller and norm of the matrices composing the Hamiltonian, the bound on the worst-case sensitivity is accurate to nearly an order of magnitude. Specifically, for the data in Figure 1, the mean of the ratio of $\beta$ to $\max\limits_{\mu}\left|\partial_{\mu}\mathsf{F}_{S}\right|$ is $23.1$.
V-B Gate Operation
@@
and optimize over the $4\mathsf{J}$-dimensional vector of control parameters $\mathbf{x}=\left(\alpha_{m1},\ldots,\alpha_{m\mathsf{J}},\omega_{m1},\ldots,\omega_{m\mathsf{J}}\right)$ for $m=1,2$. After trial and error in varying $t_{f}$ and the bound on $\set{\alpha_{mj}}$ and $\set{\omega_{mj}}$, we choose $t_{f}=2$, $|\alpha_{mj}|\leq 2$ and $0\leq\omega_{mj}\leq 10$, which generated a selection of acceptable-fidelity ($\mathsf{F}_{G}>0.99$) controllers by minimizing the objective $\mathsf{e}_{G}=1-\mathsf{F}_{G}$ following from (8) and employing the same constrained quasi-Newton optimization as before. A step size analysis similar to that of Section V-A yields $\tau=\kappa=50$ that guarantees convergence.
Figure 2: Plot of $\beta$ as given in Theorem 2 for the gate synthesis case with $|\partial_{\mu}\mathsf{F}_{S}|$ for $\mu=\set{0,1,2}$. The bound $\beta$ predicts the largest sensitivity to collective uncertainty in any of the Hamiltonian matrices to within an order of magnitude. The controllers are ordered by increasing fidelity so that the controller with index 1 yields the smallest fidelity and that with index 15 yields the greatest.
Similar to the first example, we establish an upper bound on the norm of allowable perturbations such that $\|\partial_{\mu}\Gamma^{(k)}\|\leq b=0.1287$. We then compare the bound given by $\beta=\gamma\|\mathbf{Z}\|_{1}b$ with the absolute value of the differential sensitivity computed from Section II-D. As seen in Figure 2, the bound is accurate to within an order of magnitude for all controllers with fidelity error between $10^{-2}$ and $10^{-3}$. Specifically the ratio of the bound to $\max\limits_{\mu}|\partial_{\mu}\mathsf{F}_{G}|$ has a mean of $7.73$ for this set of controllers.
VI Conclusion

### v3

@@ Abstract
{IEEEkeywords}
Quantum information and control, robust control, time-varying systems
1 Introduction
[…]
@@ 1 Introduction
We contribute to the ongoing development of robust quantum control by presenting a procedure to quantify robustness for continuously-varying, smooth controls and applicable to multi-level systems. Although piecewise constant controls have become a dominant paradigm in quantum control, such controls yield “solutions” of the Schrödinger equation that are piecewise defined and continuous but not differentiable at the switching times. Although further work is needed to explore the advantages of continuously-varying and smooth controls, there are theoretical and practical reasons why smooth controls may be preferable for certain applications. Extending our framework to evaluate robustness to such controls is therefore important. As a preview, in Section 2, we generalize the sensitivity analysis of [6] and [7] to accommodate dynamics governed by continuously varying controls. As part of this analysis, we introduce the Magnus expansion as a critical tool in the analysis of systems controlled by pulses that depart from the piecewise constant paradigm. After providing conditions for vanishing sensitivity (optimal robustness) in Section 3, we develop a bound for the sensitivity to any physically-realizable and norm-bounded uncertainty in Section 4. Finally, Section 5 presents two quantum technology-inspired case studies, based on state transfer in a spintronic network and the implementation of a quantum NOT gate.
2 Preliminaries
2.1 System Dynamics and the Magnus Expansion
@@
2.2 Approximate Solutions Given by the Magnus Expansion
@@
with $t_{k}=kh$ for $k\in\set{0,1,\ldots\tau}$. This is similar to the solution for piecewise constant controls, and the general approximation for the propagator over the interval $t_{\tau_{1}},t_{\tau_{2}})$ is
@@
Given the complexity of the higher order terms in the Magnus expansion and intractability of providing closed-form expressions for the $\Omega_{j}(t)$, typical application of the Magnus approach requires truncation of the series at a given order. Numerous methods of various orders have been derived [9]. For the numerical examples in Section 5, we adopt a fourth-order method to approximate the propagator over the time interval $t_{k-1},t_{k})$ for each $k$ from $1$ to $\tau$ as $U^{(k,k-1)}:=\exp(\Gamma^{(k)})$. Following [11, 12] the approximate dynamical generator is defined as $\Gamma^{(k)}:=$
@@
where $A^{(k)}:=A(t_{k})$, ${A}^{(k+1/2)}:=A(t_{k}+h/2)$. As this method is fourth-order, the error of the exact propagator denoted $\mathcal{U}^{(k,k-1)}$ over the time-step (k1)h(k-1)h to $kh$ compared to the truncated estimate, $\|\mathcal{U}^{(k,k-1)}-U^{(k,k-1)}\|$, is $\mathcal{O}(h^{5})$ [9, 11, 13, 14, 15].
2.3 Fidelity and Performance
@@
2.4 First-Order Sensitivity of the Performance
@@
$X_{\mu}^{(k)}=U^{(\tau,k)}\left(\partial_{\mu}\tilde{U}^{(k,k-1)}\right)U^{(k-1,0)}.$ (12)
@@ 3 Vanishing Sensitivity
[19, Th. 21.7] If the fidelity map $\mathsf{F}:U(N)\rightarrow[0,1]$ is differentiable except possibly at some isolated points, then $\forall U\in\mathsf{F}^{-1}(1)$ where the differential exists, $d_{U}\mathsf{F}=0$.
Locally charting $U(N)$ with $\mathfrak{u}(N)$, the Lie algebra of skew-Hermitian matrices, using the exponential map, and viewing the physically relevant skew-hermitian $S_{\mu}$ as providing the direction along which the sensitivity is computed, this sensitivity is $d_{{A}}\mathsf{F}\circ\mathrm{exp}(S_{\mu})$ as seen from (15).
@@
4 Bounding the Differential Sensitivity to Arbitrary Perturbations
At the other end of the spectrum, we consider upper bounds on the potential size of the differential sensitivity for a given controller for a set of allowable uncertainties or perturbations to the dynamics. To better analyze the relationship between a set of possible perturbation structures, represented by the family $\set{\partial_{\mu}\Gamma^{(k)}}$, and the resulting sensitivity, we seek an expression such that the perturbation structure encoded in $\partial_{\mu}\Gamma^{(k)}$ may be viewed as the input to a controller-specific operator that yields the resulting sensitivity to a given structure. Central to this derivation is the controller-specific mapping
@@ 4 Bounding the Differential Sensitivity to Arbitrary Perturbations
Proof 4.2.
For brevity write $\mathcal{D}^{(k)}\left(\partial_{\mu}\Gamma^{(k)}\right)$ as $X$ and $P^{(k)}z^{*}$ as $Y$. By the properties of the adjoint action of a Lie group element on its Lie algebra, $X$ is in $\mathfrak{u}(N)$ and so skew-Hermitian. Decompose $Y$ as $Y=Y_{H}+Y_{SH}$ where $Y_{H}$ is Hermitian and $Y_{SH}$ is skew-Hermitian. It then follows that $\Re\left\{\operatorname{\mathrm{Tr}}\left(XY\right)\right\}=\Re\left\{\operatorname{\mathrm{Tr}}\left(XY_{SH}\right)\right\}$ as $\Re\left\{\operatorname{\mathrm{Tr}}\left(XY_{H}\right)\right\}=0$.
@@
Proof 4.4.
The proof follows from the exposition above, which gives $\partial_{\mu}\tilde{\mathsf{F}}_{\alpha}=\gamma\sum_{k=1}^{\tau}\mathsf{z}^{(k)}\mathsf{g}_{\mu}^{(k)}$. Taking absolute values on both sides of the equation and employing the triangle equality to the sum yields $|\partial_{\mu}\mathsf{F}_{\alpha}|\leq\gamma\sum_{k=1}^{\tau}\left|\mathsf{z}^{(k)}\mathsf{g}_{\mu}^{(k)}\right|$. Applying Cauchy-Schwarz to the inner product of $\mathsf{z}^{(k)}$ with $\mathsf{g}_{\mu}^{(k)}$ for each $k$ and enforcing the bound on $\|\mathsf{g}_{\mu}^{(k)}\|\leq b$ we have $\left|\partial_{\mu}\tilde{\mathsf{F}}_{\alpha}\right|\leq\gamma\left(\sum_{k=1}^{\tau}\|\mathsf{z}^{(k)}\|\right)b=\gamma\|\mathbf{Z}\|_{1}b$.
This construction allows us to consider the controller and the input-output state information encoded in the sequence $\mathsf{z}^{(k)}$ separately from the perturbation-specific effects encoded in $\partial_{\mu}\Gamma^{(k)}$, and efficiently evaluate the worst-case sensitivity of a given controller for a collection of perturbations $\set{\partial_{\mu}\Gamma^{(k)}}_{\mu}$. Instead of bounding the sensitivity of the fidelity for a given controller with regard to certain perturbations, one can alternatively derive lower bounds on the fidelity considering a range of perturbations, an approach pursued in [21, Th. 1]. Both approaches can be seen as complementary.
5 Applications
We now apply results of Section 4 to two different quantum technology platforms subject to distinct time-varying control protocols. We do this to illustrate that the results are not dependent upon a specific application or control pulses synthesized from a specific set of basis functions.
5.1 State Transfer in Spin Ring
@@
Figure 1: Plot comparing $|\partial_{\mu}\mathsf{F}_{S}|$ for $\mu=\set{0,1,2}$ with $\beta$ from Theorem 4.3. Though computed only from the nominal system and control data, the bound has an accuracy to within slightly more than one order of magnitude. The controllers are ordered by increasing fidelity so that the controller with index 1 yields the smallest fidelity and that with index 8 yields the greatest.
Given a set of high fidelity controllers, the differential sensitivity to collective uncertainty in any of the Hamiltonian matrices is readily computed by application of (10) and (12) to (17). Based on the bound $\mathcal{B}$ used in the synthesis, we compute the maximum norm of the perturbations $\partial_{\mu}\Gamma^{(k)}$ as $b=\max\limits_{\mu,k}\|\partial_{\mu}\Gamma^{(k)}\|=0.1555$. We then test the bound $\beta$ of Theorem 4.3 after computing $\|\mathbf{Z}\|_{1}$ as derived in Section 4. Figure 1 shows the result. Without the need to explicitly compute the derivative of a matrix exponential as in (13), and based simply on knowledge of the controller and norm of the matrices composing the Hamiltonian, the bound on the worst-case sensitivity is accurate to nearly an order of magnitude. Specifically, for the data in Figure 1, the mean of the ratio of $\beta$ to $\max\limits_{\mu}\left|\partial_{\mu}\mathsf{F}_{S}\right|$ is $23.1$.
5.2 Gate Operation
@@
and optimize over the $4\mathsf{J}$-dimensional vector of control parameters $\mathbf{x}=\left(\alpha_{m1},\ldots,\alpha_{m\mathsf{J}},\omega_{m1},\ldots,\omega_{m\mathsf{J}}\right)$ for $m=1,2$. After trial and error in varying $t_{f}$ and the bound on $\set{\alpha_{mj}}$ and $\set{\omega_{mj}}$, we choose $t_{f}=2$, $|\alpha_{mj}|\leq 2$ and $0\leq\omega_{mj}\leq 10$, which generated a selection of acceptable-fidelity ($\mathsf{F}_{G}>0.99$) controllers by minimizing the objective $\mathsf{e}_{G}=1-\mathsf{F}_{G}$ following from (8) and employing the same constrained quasi-Newton optimization as before. A step size analysis similar to that of Section 5.1 yields $\tau=\kappa=50$ that guarantees convergence.
Figure 2: Plot of $\beta$ as given in Theorem 4.3 for the gate synthesis case with $|\partial_{\mu}\mathsf{F}_{S}|$ for $\mu=\set{0,1,2}$. The bound $\beta$ predicts the largest sensitivity to collective uncertainty in any of the Hamiltonian matrices to within an order of magnitude. The controllers are ordered by increasing fidelity so that the controller with index 1 yields the smallest fidelity and that with index 15 yields the greatest.
Similar to the first example, we establish an upper bound on the norm of allowable perturbations such that $\|\partial_{\mu}\Gamma^{(k)}\|\leq b=0.1287$. We then compare the bound given by $\beta=\gamma\|\mathbf{Z}\|_{1}b$ with the absolute value of the differential sensitivity computed from Section 2.4. As seen in Figure 2, the bound is accurate to within an order of magnitude for all controllers with fidelity error between $10^{-2}$ and $10^{-3}$. Specifically the ratio of the bound to $\max\limits_{\mu}|\partial_{\mu}\mathsf{F}_{G}|$ has a mean of $7.73$ for this set of controllers.
6 Conclusion